OpenStudy (anonymous):

Given the equation y − 4 =3/4 (x + 8) in point-slope form, identify the equation of the same line in standard form.

OpenStudy (jackellyn):

Distribute the 3/4 to the parenthesis then solve for y by adding 4 to both sides and simplifying.

OpenStudy (anonymous):

y-4=3/4x+6 y-10=3/4x 4y-40=3x 3x-4y+40=0

OpenStudy (anonymous):

y=3/4x+10

OpenStudy (anonymous):

3x-4y+40=0